OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (2, -2) with 0° ≤ θ < 360°.

OpenStudy (psymon):

Well, polar coordinates are (r, theta). Now we just need to know how we change x and y into r and theta. To start with r^2 = x^2 + y^2 tan(theta) = y/x So this is how we start off. Think you would be able to get r or theta using those two conversions?

OpenStudy (anonymous):

I'll try, one sec

OpenStudy (psymon):

No prob :3

OpenStudy (anonymous):

Is r 2sqrt2 and theta 1?

OpenStudy (psymon):

tan(theta) = -1 actually. The radius is correct, though.

OpenStudy (anonymous):

oh,ok

OpenStudy (psymon):

Now, what quadrant would (2,-2) be in?

OpenStudy (anonymous):

the 4th

OpenStudy (psymon):

Right. So in the 4th quadrant, where does tan(theta) = -1?

OpenStudy (anonymous):

3pi/2?

OpenStudy (psymon):

That is where tangent is undefined actually.

OpenStudy (psymon):

Tangent is sin/cos. Theonly way we could end up with a -1 is if sin and cos are the same value, one being negative. So there are only 4 angles where sin and cos are the same, one in each quadrant.

OpenStudy (anonymous):

7pi/4?

OpenStudy (psymon):

Bingo. So our first pair would be (2rt2, 7pi/4). Now, what this is actually saying is that if we face in the direction of 7pi/4, we walk forward 2rt2 unitsto get to our coordinate. Now, there are two ways we can get to our coordinate. We can face in that direction and walk forward OR we can face in the opposite direction and walk backwards. So what angle would be in the OPPOSITE direction?

OpenStudy (anonymous):

Um, would it be 5pi/4?

OpenStudy (psymon):

Nope. We would want to face 180 degrees in the opposite direction. Basically, we want to subtract pi from our angle, that would be how we face the opposite direction.

OpenStudy (anonymous):

3pi/4?

OpenStudy (psymon):

Yep :3 So our angle is 3pi/4. Now the radius is the same, but when we walk backwards, our radius is negative, making our 2nd pair (-2rt2, 3pi/4)

OpenStudy (anonymous):

thank you so much! I actually understand everything better now! :)

OpenStudy (psymon):

cool ^_^ glad to hear.